OpenStudy (anonymous):

Solve each inequality. Graph the solution set and answer in interval notation. Terrible at math walk me through the steps Please Help!!!!!!!! -17 less then 3/2c -5 less then 1

OpenStudy (anonymous):

\[-17< \frac{ 3 }{ 2 }c-5<1\]

OpenStudy (campbell_st):

add 5 to both sides of the inequality \[-12 < \frac{3}{2} c < \] multiply by 2 then divide by 3 this will give you the interval...

OpenStudy (anonymous):

add 5 to all the sides u will get -12<3/2 c<6 now multiply throughout by 2 -24<3c<12 now divide throughout by 3 -8<c<4

OpenStudy (campbell_st):

oops should read \[-12 < \frac{3}{2} c < 6\]